OpenStudy (anonymous):

There are three red marbles, four white marbles, five yellow marbles, and six black marbles in a bag. What is the probability of drawing two red marbles, if you don't replace the marble after you draw the first one? A.) 1/51 B.) 1/6 C.) 1/36 D.) 1/54

OpenStudy (texaschic101):

total marbles : 3 + 4 + 5 + 6 = 18 red marble on first try = 3/18 = 1/6 red marble on second try = 2/17 now multiply......2/17 x 1/6 = 2/102 = 1/51 you might want to get a second opinion on this....just to be sure. Probability's are not my best subject.

OpenStudy (anonymous):

Good enough! Thanks!

OpenStudy (texaschic101):

I hope it is correct :/

OpenStudy (anonymous):

It seems like it will be. I mean like the work seems to be correct